Bob is 6’, weighs 225 lbs, and is 40 years old. His waist measurement is 42 inches and his hip measurement is 36 inches.

Food Recall: Bkf=2 donuts, 3 cups coffee with creamer

            Lunch: fast food—cheeseburger, large fries, large coke

            Dinner: Fried chicken, mashed potatoes and gravy, green beans, canned

               peaches

            Snack: ice cream

4. What is his BMI?

5. How is his BMI classified?

6. What are his total calorie needs? (REE x activity factor x 1.1)

REE = 1.0 kcal/kg/hr                                         Light: 1.3-1.45 (30-45%)

Thermic effect of food (5% to 10%) (1.1)

7. What type of body fat distribution do you think Bob has?

Bob is interested in weight loss.

8. What should he look for in a weight loss program?

9. How many calories should he consume to help with weight loss?

10. What is a good percentage of total weight that is a realistic weight-loss target in six months?

11. Make food suggestions based on Bob’s food recall to help him lower his calorie

      intake.
Answer #1

4. What is his BMI?

BMI formula = (weight in kg) / [height in (m)2]

Convert weight in kg.

Bob weight = 225 pounds

1kg = 2.2 pounds

Weight in kg = 225/ 2.2

weight in kg = 102.2

Convert height into metre

Bob height = 6feet

divide by 3.281

Height = 6 / 3.281

Height = 1.82m

BMI = 102.2 / (1.82)2

BMI = 102.2 / 3.31

BMI = 30.8kg/m2

5. How is his BMI classified?

His BMI is classified as obese Class I. BMI between 30 - 35 is considered as an obese class I.

6. What are his total calorie needs?

Total calorie needs = BMR x Light exercise

Calculate BMR using the equation

BMR = 66.4730 + (13.7516 x weight in kg) + (5.0033 x height in cm) – (6.7550 x age in years)

BMR = 66.4730 +(13.7516 x 102.2) + (5.0033 x 182.8) - (6.7550 x 40)

= 66.4730 + 1405.4+ 914.6 - 270.2

= 66.4730 + 1405.4 + 644.4

BMR = 2116.2

Total calorie needs = BMR x 1.3

= 2116.2 X 1.3

Total calorie needs = 2751

7. What type of body fat distribution do you think Bob has?

Bob has 30% of body fat which is obese for his age. He has muscle body mass fat distribution.